OpenStudy (anonymous):

What is the 7th term of the geometric sequence -1, 4, -16

OpenStudy (cwrw238):

7th term = ar^6 a = -1 and r = 4/-1

OpenStudy (anonymous):

\[a _{n} = a _{1}r ^{n - 1}\] \[a _{n} = -1(-4) ^{7 - 1}\] \[a _{n} = -1(-4) ^{6}\] \[a _{n} = -1(4096)\] \[a _{n} = -4096\]

OpenStudy (anonymous):

After the first equation, it should be a_7

OpenStudy (anonymous):

Not a_n

OpenStudy (ujjwal):

the \(n^{th}\) term of Geometric sequence is given by a\(r^{n-1}\) where a is first term. r is common ratio.